What is the probability of getting a 2 or a 5 when a dice is rolled?


1
Expert's answer
2020-08-18T16:53:14-0400

The probability of getting 2 is 1/6 ( dice has 6 sides, we need one of them).

Similary the probability of getting 5 is 1/6. Further the probability of getting 2 or 5 is computed with a help of addition of probabilities: 1/6+1/6= 2/6=1/3.